OpenStudy (anonymous):

2log7 ^6 - log7x = 1/2log7 ^16 7 is the subscript for each of them

OpenStudy (kinggeorge):

First combine it all into a single log function. Once you've done that you can just remove the logs and set the interiors equal to each other. Is your function\[2\;\log_7{(6)}-\log_7{(x)} = 1/2\; \log_7{\;(16)} ?\]

OpenStudy (kinggeorge):

If you wanted to combine it into a single log function you would have to do the following.\[2\;\log_7(6)−\log_7(x)=1/2\;\log_7\;(16)\]\[\log_7(6^2)−\log_7(x)=\log_7\;(16^{1/2}) \]\[\log_7\left({6^2 \over x}\right)=\log_7\;(16^{1/2})\]From there you can just remove the log function and solve for x.